# Quantum Rationale-Aware Graph Contrastive Learning for Jet Discrimination

### v1

@@ Abstract
In high-energy physics, particle jet tagging plays a pivotal role in distinguishing quark from gluon jets using data from collider experiments. While graph-based deep learning methods have advanced this task beyond traditional feature-engineered approaches, the complex data structure and limited labeled samples present ongoing challenges. However, existing contrastive learning (CL) frameworks struggle to leverage rationale-aware augmentations effectively, often lacking supervision signals that guide the extraction of salient features and facing computational efficiency issues such as high parameter counts. In this study, we demonstrate that integrating a quantum rationale generator (QRG) within our proposed Quantum Rationale-aware Graph Contrastive Learning (QRGCL) framework significantly enhances jet discrimination performance, reducing reliance on labeled data and capturing discriminative features. Evaluated on the quark-gluon jet dataset, QRGCL achieves an AUC score of 77.53% while maintaining a compact architecture of only 45 QRG parameters, outperforming classical, quantum, and hybrid GCL and GNN benchmarks. These results highlight QRGCL’s potential to advance jet tagging and other complex classification tasks in high-energy physics, where computational efficiency and feature extraction limitations persist.
Index Terms:
quantum machine learning, graph neural network, contrastive learning, high-energy physics, jet classification
I Introduction
@@
Particle jet tagging, a fundamental task in high-energy physics, aims to identify the originating parton-level particles by analyzing collision byproducts at the Large Hadron Collider (LHC). While traditional approaches have relied on manually engineered features, modern deep learning methods offer promising alternatives for processing vast amounts of collision data [1]. The representation of jets as collections of constituent particles has emerged as a more natural and flexible approach compared to image-based methods, allowing for the incorporation of arbitrary particle features [2]. However, the challenge of limited labeled data in particle physics necessitates innovative solutions beyond purely supervised learning approaches. Self-supervised pretraining followed by supervised fine-tuning has shown particular promise in this domain. Self-supervised contrastive learning (CL) [3] has gained significant attention in the field of graph neural networks (GNNs) [4], leading to the development of graph CL (GCL). This approach involves pre-training a GNN on large datasets without manually curated annotations, facilitating effective fine-tuning for subsequent tasks [5].
@@
A review of existing GCL approaches reveals a common framework that combines two primary modules: (1) graph augmentation, which generates diverse views of anchor graphs through techniques like node dropping, edge perturbation, and attribute masking, and (2) CL, which maximizes agreement between augmented views of the same anchor while minimizing agreement between different anchors. However, these methods face inherent challenges due to the complexity of graph structures, where random augmentations may obscure critical features, potentially misguiding the CL process. In response to these challenges, recent studies have shifted focus towards understanding the invariance properties of GCL. The necessity for augmentations was emphasized to maintain semantic integrity, arguing that high-performing GCL frameworks should enhance instance discrimination without compromising the intrinsic semantics of the anchor graphs. Building on this foundation, invariant rationale discovery (IRD) techniques [3] were proposed, aligning closely with the objectives of GCL. These techniques highlight the importance of identifying critical features that inform predictions effectively.
@@
Despite these advancements, gaps remain in effectively leveraging rationales for augmentation. Existing frameworks often lack supervision signals, hindering their ability to reveal and utilize the most salient features effectively. As the amount of available data increases and the computational cost of deep learning networks rises, significant computing resources will be required to execute these algorithms efficiently. To handle computational complexity, we see a move from classical networks that process bits to quantum networks that use qubits. Quantum networks leverage unique properties such as superposition and entanglement, allowing them to represent $2^{n}$ characteristics from $n$ two-dimensional complex vectors. In contrast to classical networks, which scale linearly in expressiveness, quantum networks show exponential growth as the sample size $n$ increases [6, 7]. Recognizing the pivotal role of the rationale generator in the GCL framework, we propose enhancing this component with a quantum-based subroutine to evaluate its potential for performance improvement. Our proposed Quantum Rationale-aware Graph Contrastive Learning (QRGCL) addresses these limitations by integrating a quantum rationale generator (QRG) that autonomously identifies and reveals salient features within the graphs, ensuring that the generated augmentations retain their discriminative power. Our approach incorporates the state-of-the-art ParticleNet [2] GNN encoder network followed by a projection head, culminating in a multi-layer perceptron (MLP) classifier. Experimental results on the quark-gluon jet tagging dataset showcase the superiority of our approach, proving its capacity to capture distinguishing semantic nodes and significantly outperforming current state-of-the-art GCL and GNN methods (classical, quantum, and hybrid).
The paper is structured as follows: Section II reviews related work, while Section III describes our dataset, the proposed QRGCL model, and the benchmark models. Section IV discusses the experimental setup, and Section V presents the results and discussion. Finally, Section VI summarizes the key findings and suggests directions for future research.
II Background
II-A Contrastive Representation Learning
Contrastive representation learning [8, 9, 10, 11] is an effective framework for extracting meaningful representations from high-dimensional data by mapping it into a lower-dimensional space. CL uses a self-supervised approach to differentiate between positive pairs (similar data) and negative pairs (dissimilar data), particularly when labeled data is scarce. The framework consists of three components:
  1. 1.
  2. 2.
  3. 3.
II-B Quantum Contrastive Learning (QCL)
Recent studies have begun to explore how quantum computing can enhance traditional CL frameworks. Quantum systems can potentially offer superior computational capabilities, allowing for more complex feature extraction and representation learning. One study [12] proposed a hybrid quantum-classical model for self-supervised learning, showing that small-scale QNNs could effectively enhance visual representation learning. By training quantum and classical networks together to align augmented image views, they achieved higher test accuracy in image classification than a classical model alone, even with limited quantum sampling. Another approach [9] integrates Supervised CL (SCL) with Variational Quantum Circuits (VQC) and incorporates Principal Component Analysis (PCA) for effective dimensionality reduction. This method addresses the limitations posed by scarce training data and showcases potential in medical image analysis. Experimental results reveal that this model achieves impressive accuracy across various medical imaging datasets, particularly with a minimal number of qubits (2 qubits), underscoring the benefits of quantum computing. A different research effort presents Q-SupCon [10], a fully quantum-enhanced Supervised CL (SCL) model tailored to tackle issues related to data scarcity. Experiments demonstrate that this model yields significant accuracy in image classification tasks, even with very limited labeled datasets. Its robust performance on actual quantum devices illustrates its adaptability in scenarios with constrained data availability. Furthermore, a quantum-enhanced self-supervised CL framework has been proposed for effective mental health monitoring [11]. This framework leverages a quantum-enhanced Long Short-Term Memory (LSTM) encoder to enhance representation learning for time series data through CL. The results indicate that this model significantly outperforms traditional self-supervised learning approaches, achieving high F1 scores across multiple datasets. To take advantage of QCL, as evident in these studies, we attempted to replace the classical rationale generator with a VQC in our proposed framework.
II-C Rationale-Aware Graph Contrastive Learning (RGCL) Concept
RGCL [3] represents a self-supervised CL approach that overcomes several limitations common to traditional graph CL (GCL) frameworks. Standard GCL methods often suffer from challenges such as augmentation strategies that can inadvertently alter or remove critical graph structure and semantics, and attempts to preserve graph-specific domain knowledge sometimes result in overfitting, limiting the model’s adaptability to diverse, unseen data. RGCL addresses these issues by focusing on the concept of rationale learning, where the essential, discriminative information for graph classification is typically concentrated within a subset of nodes or edges in the graph. In RGCL, this discriminative subset, or rationale, is identified and emphasized during training, allowing the model to prioritize meaningful patterns while minimizing reliance on irrelevant features. In RGCL, specialized neural networks, known as the rationale generator (RG), are used to assign importance scores to each node. This generator evaluates each node’s contribution to the graph’s overall representation. The higher-score nodes form the rationale subset, while the remaining nodes comprise the complement subset. The rationale subset undergoes targeted augmentations during training, capturing the core discriminative features essential for downstream tasks. In contrast, the complement subset is augmented to encourage the exploration of less critical correlations, thus avoiding overfitting and improving generalization. RGCL pipeline leverages these dual views—rationale and complement—to guide the encoder network in learning a balanced feature space. By focusing on rationale views, the model learns robust, task-relevant features, while the complement views prevent it from becoming overly sensitive to spurious relationships, fostering a more generalized understanding.
@@ II-C Rationale-Aware Graph Contrastive Learning (RGCL) Concept
Figure 1: QRGCL framework. The quantum rationale generator identifies a discriminative subset of nodes in the original graph. The rationale generator shared GNN-encoder, and projection head is jointly optimized by minimizing the combined loss.
III Methodology
III-A Dataset and Preprocessing
III-A1 High-Energy Physics Dataset
@@
III-A2 Graph Representation of Jets
@@
III-A3 Feature Engineering
$m_{\alpha,T}^{(i)}=\sqrt{m_{\alpha}^{(i)^{2}}+p_{\alpha,T}^{(i)^{2}}}$ (1)
$E_{\alpha}^{(i)}=m_{\alpha,T}^{(i)}\text{cosh}y_{\alpha}^{(i)}$ (2)
$\vec{p}_{\alpha}=\sum_{i}\vec{p}_{\alpha}^{(i)},$ (5)
with the transverse momentum of the jet ($p_{T,\alpha}$) calculated as:
III-A4 Edge Construction and Attributes
@@
III-A5 Graph-Based Augmentation and Contrastive Learning
Next, the preprocessed graph data creates pairs or “views” as input for our CL framework. In CL, pairs of similar and dissimilar views are generated to help the model learn discriminative representations. Positive views are created by taking a graph and generating an augmented version of it, such as applying transformations like node dropping, edge perturbation, or feature masking. For instance, an augmented view of a quark jet remains labeled as similar (1), while a dissimilar pair may consist of a quark jet and a gluon jet, labeled as 0. The differentiation between positive and negative pairs is established solely through the loss function, with the model lacking an inherent understanding of the concept of ‘view’. The loss function guides the model toward clustering similar samples in proximity. Figure 3b shows positive and negative pairs created for our CL process. The distorting jets method was applied to shift the positions of the jet constituents independently, with shifts drawn from a normal distribution. The shift is applied to each constituent’s $\eta$ and $\psi$ values, scaled by their $p_{T}$, ensuring that lower $p_{T}$ particles experience more significant shifts. The collinear fill technique added collinear splittings to the jets, filling zero-padded entries by splitting existing particles. A random proportion is applied for each selected particle to create two new particles that share the original momentum and position information.
III-A6 Enforcing Infrared and Collinear Safety
III-A7 Data Splitting
@@ III-B Proposed QRGCL
III-B1 Quantum Rationale Generator (QRG)
The QRGCL model substitutes its classical RG (CRG) with a quantum RG (QRG). This component is crucial in generating augmented graph representations by assigning significance scores to each node. The QRG is built using a 7-qubit parameterized quantum circuit (PQC), where each qubit represents a node in the graph.
@@ III-B1 Quantum Rationale Generator (QRG)
Figure 4: QRG circuit of the proposed QRGCL. The circuit operates on seven qubits, with each qubit corresponding to a node in the graph. The top portion shows the data encoding stage, where each qubit is initialized using a $H$ gate followed by $RX$-based angle encoding node features. $CRZ$ gates encode edge relationships between qubits. The bottom portion includes parameterized rotations ($RX$, $RY$, $RZ$) for adaptable representations and entanglement layers using $SWAP$ gates. Measurement results are obtained on a computational basis, with classical registers collecting the output.
@@
III-B2 Encoder Network
The input to the encoder is a matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$, where $n$ represents the number of particles and $d$ the feature dimension (e.g., momentum, energy). ParticleNet constructs a k-nearest neighbor (k-NN) graph, connecting each particle to its $k$ closest neighbors in feature space. The EdgeConv block begins by computing the $k$ nearest neighbors using the particles’ spatial coordinates. Edge features are constructed based on the feature vectors of these neighbors. The core operation of EdgeConv is a 3-layer multi-layer perceptron (MLP), where each layer consists of a linear transformation, batch normalization, and a ReLU activation. To improve information flow and avoid vanishing gradients, a shortcut connection inspired by general ResNet architecture runs parallel to the EdgeConv operation, allowing input features to bypass the convolution layers. The two main hyperparameters of each EdgeConv block are $k$, the number of nearest neighbors, and $C=(C_{1},C_{2},C_{3})$, the number of units in each MLP layer.
III-B3 Projection Head
@@
Our architecture incorporates a projection head that effectively maps high-dimensional embeddings from the encoder into compact, CL-optimized latent space. This distinction allows the encoder to learn features suited for downstream tasks while the projection head focuses on maximizing contrastive effectiveness. This projection head employs an initial linear transformation to reduce the dimensionality of the input embedding (128-dimensional) to a latent space of the same dimensionality, followed by a ReLU activation and a final linear layer. This layered design allows the projection head to align representations tightly with the CL objective by enforcing mutual information between anchor and rationale representations in the latent space.
III-B4 Quantum-Enhanced Contrastive Loss
@@
QRGCL model uses a carefully designed loss function that integrates multiple elements—InfoNCE, alignment, uniformity, rationale-aware loss (RA loss), and contrastive pair loss (CP loss)—to optimize the learning of quantum-enhanced embeddings.
InfoNCE Loss: The InfoNCE loss focuses on distinguishing positive pairs from negative pairs. It is defined as:
$\mathcal{L}_{\text{InfoNCE}}=-\log\left(\frac{e^{(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{2})/T)}}{\sum_{j=1}^{N}\mathbb{1}_{j\neq i}e^{(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{j})/T)}}\right)$ (14)
where $\mathbf{z}_{i}$ and $\mathbf{z}_{2}$ are embeddings of a positive pair, $T$ is the temperature parameter, and $N$ is the batch size. The temperature parameter helps control the sharpness of the similarity scores, making the model more or less sensitive to differences between pairs.
RA Loss: The RA loss is designed to focus on the “rationale” of the graph—subsets of nodes that contain critical classification information. This is calculated as:
$\mathcal{L}_{\text{RA}}=-\log\left(\frac{e^{(\text{sim}(\mathbf{z}_{1},\mathbf{z}_{2})/T)}}{\sum_{j=1}^{N}e^{(\text{sim}(\mathbf{z}_{1},\mathbf{z}_{j})/T)}-e^{(\text{sim}(\mathbf{z}_{1},\mathbf{z}_{2})/T)}}\right)$ (15)
The RA loss ensures that the model concentrates on similarities between positive pairs while minimizing the influence of other relationships.
CP Loss: The CP loss encourages the model to distinguish between positive pairs and a complement set (pairs involving nodes not deemed crucial for the classification task):
$\mathcal{L}_{\text{CP}}=-\log\left(\frac{e^{(\text{sim}(\mathbf{z}_{1},\mathbf{z}_{2})/T)}}{\sum_{j=1}^{N}e^{(\text{sim}(\mathbf{z}_{1},\mathbf{z}_{3})/T)}+e^{(\text{sim}(\mathbf{z}_{1},\mathbf{z}_{2})/T)}}\right)$ (16)
where $\mathbf{z}_{3}$ represents embeddings from the complement set. This loss helps the model differentiate genuine relationships between similar samples from spurious correlations by learning from pairs with varying significance.
Alignment Loss: It measures the closeness of normalized embeddings of positive pairs ($p_{pos}$):
$\mathcal{L}_{\text{align}}\triangleq\mathbb{E}_{(x,y)\sim p_{pos}}\left[\|\text{normalize}(\mathbf{z}_{1})-\text{normalize}(\mathbf{z}_{2})\|^{2}\right]$ (17)
We used quantum state fidelity as a distance metric, replacing the traditional $L^{2}$ distance typically used in classical CL. Quantum fidelity measures the closeness between two quantum states. Thus, alignment loss becomes:
$\mathcal{L}_{\text{align}}\triangleq\mathbb{E}_{(x,y)\sim p_{pos}}[1-\mathcal{F}(\textbf{z}_{1},\textbf{z}_{2})]=1-\left(\text{Tr}\left(\sqrt{\sqrt{\textbf{z}_{1}}\textbf{z}_{2}\sqrt{\textbf{z}_{1}}}\right)\right)^{2}$ (18)
where $\rho_{1}$ and $\rho_{2}$ are density matrices of the quantum states. This metric replaces the usual Euclidean distance in the alignment term of the loss function.
Uniformity Loss: It ensures that embeddings are uniformly distributed over the hypersphere, avoiding representation collapse:
$\mathcal{L}_{\text{uniform}}\triangleq\log\mathbb{E}_{(x,y)\sim p_{data}}\left[e^{(-2\,\text{pdist}(\mathbf{z}))}\right]$ (19)
where pdist calculates the pairwise distances between embeddings. This helps maintain representation diversity, allowing the model to capture more nuanced differences.
The overall loss for the QRGCL model is a weighted combination of the InfoNCE, RA, and CP loss, with optional contributions from the alignment and uniformity terms:
where $\lambda$, $\alpha$, $\beta$, and $\delta$ are hyperparameters that control the contributions of each loss component. During experimentation, the uniformity term hindered learning, so it can be omitted when necessary.
III-B5 Classifier Head

### v2

@@ Abstract
In high-energy physics, particle jet tagging plays a pivotal role in distinguishing quark from gluon jets using data from collider experiments. While graph-based deep learning methods have advanced this task beyond traditional feature-engineered approaches, the complex data structure and limited labeled samples present ongoing challenges. However, existing contrastive learning (CL) frameworks struggle to leverage rationale-aware augmentations effectively, often lacking supervision signals that guide the extraction of salient features and facing computational efficiency issues such as high parameter counts. In this study, we demonstrate that integrating a quantum rationale generator (QRG) within our proposed Quantum Rationale-aware Graph Contrastive Learning (QRGCL) framework significantly enhances jet discrimination performance, reducing reliance on labeled data and capturing discriminative features. Evaluated on the quark-gluon jet dataset, QRGCL achieves an AUC score of 77.53% while maintaining a compact architecture of only 45 QRG parameters, outperforming classical, quantum, and hybrid GCL and GNN benchmarks. These results highlight QRGCL’s potential to advance jet tagging and other complex classification tasks in high-energy physics, where computational efficiency and feature extraction limitations persist. The source code for QRGCL is available at https://anonymous.4open.science/r/QRGCL-1951.
1 Introduction
@@
Particle jet tagging, a fundamental task in high-energy physics, aims to identify the originating parton-level particles by analyzing collision byproducts at the Large Hadron Collider (LHC). While traditional approaches have relied on manually engineered features, modern deep learning methods offer promising alternatives for processing vast amounts of collision data [1]. The representation of jets as collections of constituent particles has emerged as a more natural and flexible approach compared to image-based methods, allowing for the incorporation of arbitrary particle features [2, 3]. However, the challenge of limited labeled data in particle physics necessitates innovative solutions beyond purely supervised learning approaches. Self-supervised pretraining followed by supervised fine-tuning has shown particular promise in this domain. Self-supervised contrastive learning (CL) [4, 5, 6, 7] has gained significant attention in the field of graph neural networks (GNNs) [8, 9], leading to the development of graph CL (GCL). This approach involves pre-training a GNN on large datasets without manually curated annotations, facilitating effective fine-tuning for subsequent tasks [10].
@@
A review of existing GCL approaches reveals a common framework that combines two primary modules: (1) graph augmentation, which generates diverse views of anchor graphs through techniques like node dropping, edge perturbation, and attribute masking, and (2) CL, which maximizes agreement between augmented views of the same anchor while minimizing agreement between different anchors. However, these methods face inherent challenges due to the complexity of graph structures, where random augmentations may obscure critical features, potentially misguiding the CL process. In response to these challenges, recent studies have shifted focus towards understanding the invariance properties [11, 12] of GCL. The necessity for augmentations was emphasized to maintain semantic integrity, arguing that high-performing GCL frameworks should enhance instance discrimination without compromising the intrinsic semantics of the anchor graphs. Building on this foundation, invariant rationale discovery (IRD) techniques [7, 13, 14] were proposed, aligning closely with the objectives of GCL. These techniques highlight the importance of identifying critical features that inform predictions effectively.
@@
Despite these advancements, gaps remain in effectively leveraging rationales for augmentation. Existing frameworks often lack supervision signals, hindering their ability to reveal and utilize the most salient features effectively. As the amount of available data increases and the computational cost of deep learning networks rises, significant computing resources will be required to execute these algorithms efficiently [15]. To handle computational complexity, we see a move from classical networks that process bits to quantum networks that use qubits [16, 17]. Quantum networks leverage unique properties such as superposition and entanglement, allowing them to represent $2^{n}$ characteristics from $n$ two-dimensional complex vectors. In contrast to classical networks, which scale linearly in expressiveness, quantum networks show exponential growth as the sample size $n$ increases [18, 19]. Recognizing the pivotal role of the rationale generator in the GCL framework, we propose enhancing this component with a quantum-based subroutine to evaluate its potential for performance improvement. Our proposed Quantum Rationale-aware Graph Contrastive Learning (QRGCL) addresses these limitations by integrating a quantum rationale generator (QRG) that autonomously identifies and reveals salient features within the graphs, ensuring that the generated augmentations retain their discriminative power. Our approach incorporates the state-of-the-art ParticleNet [2] GNN encoder network, followed by a projection head, culminating in a multi-layer perceptron (MLP) classifier. Experimental results on the quark-gluon jet tagging dataset showcase the superiority of our approach, proving its capacity to capture distinguishing semantic nodes and significantly outperforming current state-of-the-art GCL and GNN methods (classical, quantum, and hybrid).
Our main contributions are: (i) We propose a novel hybrid quantum-classical framework, Quantum Rationale-aware Graph Contrastive Learning (QRGCL), that integrates a quantum rationale generator to identify salient substructures in graph-structured particle physics data for enhanced contrastive learning; (ii) We design a parameter-efficient QRG based on a 7-qubit variational quantum circuit, enabling salient feature extraction with only 45 trainable parameters; (iii) We introduce a new quantum-enhanced contrastive loss that incorporates rationale-aware, contrastive pairs, and alignment losses, with quantum fidelity as a distance metric; (iv) We conduct extensive experiments on the quark-gluon jet tagging dataset, showing that QRGCL significantly outperforms classical, quantum, and hybrid benchmarks in terms of AUC, while maintaining a compact and computationally efficient architecture.
2 Background
2.1 Contrastive Representation Learning
CL [20, 21, 22, 23] is a self-supervised framework that learns discriminative representations by pulling together embeddings of similar samples (positive pairs) and pushing apart dissimilar ones (negative pairs). It typically involves data augmentation, an encoder, and a projection head. We detail these components and their role in jet physics in Appendix A.1.
2.2 Quantum Contrastive Learning (QCL)
Recent work has explored quantum enhancements to CL. Models such as Q-SupCon [22], quantum-PCA-based CL [21], and QCLR [23] demonstrate that quantum circuits can improve CL performance in data-scarce settings. Inspired by these findings, we replace the classical rationale generator with a VQC in our proposed framework. Further discussions are provided in Appendix A.2.
2.3 Rationale-Aware Graph Contrastive Learning (RGCL)
RGCL [7] addresses the limitations of traditional GCL by focusing on discriminative substructures, known as rationales. It uses a rationale generator to separate salient and non-salient graph components, which are then processed through distinct augmentations to promote robust and generalizable representations. We provide detailed illustrations of the RGCL pipeline in Appendix A.3.
@@ 2.3 Rationale-Aware Graph Contrastive Learning (RGCL)
Figure 1: QRGCL framework. The quantum rationale generator identifies a discriminative subset of nodes in the original graph. The rationale generator shared GNN-encoder and projection head is jointly optimized by minimizing the combined loss.
3 Methodology
3.1 Dataset and Preprocessing
3.1.1 High-Energy Physics Dataset
@@
3.1.2 Graph Representation of Jets
@@
3.1.3 Feature Engineering
Additional kinematic variables are derived from the original features ($p_{T,\alpha}^{(i)},y_{\alpha}^{(i)},\psi_{\alpha}^{(i)}$) using the ‘Particle‘ package to enhance the model’s ability to learn from the data. These engineered features include transverse mass, energy, and Cartesian momentum components, providing a more complete description of each particle’s dynamics. More details about these engineered features are shown in Appendix B.1.
3.1.4 Edge Construction and Attributes
@@
3.1.5 Graph-Based Augmentation and Contrastive Learning
We applied graph-based CL techniques to enhance the discriminative capability of our models by generating augmented graph views. The augmentation strategies included node dropping, edge perturbation, feature masking, and jet-specific transformations. Details of the augmentation pipeline, view pairing, and the construction of positive and negative pairs are provided in Appendix B.2.
3.1.6 Enforcing Infrared and Collinear Safety
To ensure compliance with infrared and collinear (IRC) safety, we adopt perturbation and regularization techniques inspired by the principles of QCD, as outlined by Dillon et al. [26]. Further theoretical and implementation details are provided in Appendix B.2.1.
3.1.7 Data Splitting
@@ 3.2 Proposed QRGCL
3.2.1 Quantum Rationale Generator (QRG)
The QRGCL model substitutes its classical RG (CRG) [7] with a quantum RG (QRG). This component is crucial in generating augmented graph representations by assigning significance scores to each node. The QRG is built using a 7-qubit parameterized quantum circuit (PQC), where each qubit represents a node in the graph.
@@ 3.2.1 Quantum Rationale Generator (QRG)
Figure 4: QRG circuit of the proposed QRGCL. The circuit operates on seven qubits, with each qubit corresponding to a node in the graph. The top portion shows the data encoding stage, where each qubit is initialized using an $H$ gate followed by $RX$-based angle encoding node features. $CRZ$ gates encode edge relationships between qubits. The bottom portion includes parameterized rotations ($RX$, $RY$, $RZ$) for adaptable representations and entanglement layers using $SWAP$ gates. Measurement results are obtained on a computational basis, with classical registers collecting the output.
@@
3.2.2 Encoder Network
We used the ParticleNet [2] model as our encoder to convert the augmented views of input particle features into low-dimensional embeddings. ParticleNet is a graph-based neural network optimized for jet tagging, leveraging dynamic graph convolutional neural networks (DGCNN) to process unordered sets of particles, treating jets as particle clouds. More details about ParticleNet are shown in Appendix C.1.
3.2.3 Projection Head
@@
Our architecture incorporates a projection head that effectively maps high-dimensional embeddings from the encoder into a compact, CL-optimized latent space. This distinction allows the encoder to learn features suited for downstream tasks while the projection head focuses on maximizing contrastive effectiveness. This projection head employs an initial linear transformation to reduce the dimensionality of the input embedding (128-dimensional) to a latent space of the same dimensionality, followed by a ReLU activation and a final linear layer. This layered design allows the projection head to align representations tightly with the CL objective by enforcing mutual information between anchor and rationale representations in the latent space.
3.2.4 Quantum-Enhanced Contrastive Loss
@@
QRGCL model uses a carefully designed loss function that integrates multiple elements: InfoNCE, alignment, uniformity, rationale-aware loss (RA loss), and contrastive pair loss (CP loss), to optimize the learning of quantum-enhanced embeddings. More details about these losses are shown in Appendix C.2.
3.2.5 Classifier Head